OpenStudy (anonymous):

*correction The vertices of the hyperbola 25y^(2)-9x^(2)-200y+36x+139=0 are ______

OpenStudy (amistre64):

(25y^2-200y+___) + (-9x^2 +36x +___) = -139 25(y^2-8y+(16)) -9 (x^2 -4x +(4)) = -139-36+400 25(y-4)^2 - 9(x-2)^2 = 171 25(y-4)^2 9(x-2)^2 --------- - -------- = 1 171 171 accordogng to this setup; the verts are at (2, 4 +-sqrt(171)/25)

OpenStudy (amistre64):

(2, 4 +-sqrt(171)/5) that is; if i did it right